OpenStudy (anonymous):

Several systems of equations are given below. System 1 y = 6x – 1.5 y = –6x + 1.5 System 2 x + 3y = –6 2x + 6y = 3 System 3 2x –y = 5 6x – 3y = 15 Which system of equations is consistent-independent? How many solutions will the system of equations have? Expain your answers. B:Which system of equations is consistent-dependent? How many solutions will the system of equations have? Expain your answers. C:Which system of equations is inconsistent-independent? How many solutions will the system of equations have? Expain your answers.

OpenStudy (anonymous):

Just graph each group of systems at a time and graph online

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

There are some things so savvy 1. If there is a solution to the equations, then they are consistent. 2. If there is only one solution, they are independent. 3. If they are the same equation, then they are consistent and dependent. 4. if they are parallel they are inconsistent. Now you have to solve the equation; you have to set one equal to another and solve for x Looks like y = 6x – 1.5 y = –6x + 1.5 Now for the second line in the equation it would be the same as the first; multiplied by -1. \[6x - 1.5 = -6x + 1.5\] \[12x = 3\] \[x = \frac{3}{12}\] \[x = \frac{1}{4}\] this is the system of equations is consistent and independent.